OpenStudy (anonymous):

Alexus deposited $1900 into a savings account that pays a simple annual interest rate of 1.8%. How much interest will she earn after 9 months?

OpenStudy (anonymous):

$25.65

OpenStudy (anonymous):

$1900*1.8=34.20/12=2.85*9=25.65

OpenStudy (anonymous):

sry subtract 34.20-25.65=8.55